OpenStudy (anonymous):

1. 7 1/2 of what # is 4 1/2? and 2. Evaluate: 3^2-2^4+/- "divide sign" 25

OpenStudy (campbell_st):

75x = 4.5 x = 4.5/7.5 x = 3/5

OpenStudy (campbell_st):

so its (9 - 16)/25 = - 7/25 or 9 - 16/25... 8 9/25

OpenStudy (anonymous):

Thank you so much! :)